OpenStudy (anonymous):

Around the year 600 B.C., Babylon, the capital of Babylonia, was the largest city in the world. It covered an area of 2500 acres. If one acre is approximately equivalent to 1.56 x 10 to the negative 3rd power miles squared, find the area of Babylon in square miles.

OpenStudy (anonymous):

please show me how to do it too.

OpenStudy (xishem):

\[\frac{2500acres}{1}*\frac{1.56*10^{-3} mi^2}{1acre}=3.90mi^{2}\]

OpenStudy (anonymous):

Thank you!